# DRT-o1: Optimized Deep Reasoning Translation via Long Chain-of-Thought

### v1

@@ Abstract
Recently, O1-like models have emerged as representative examples, illustrating the effectiveness of long chain-of-thought (CoT) in reasoning tasks such as math and coding tasks. In this paper, we introduce DRT-o1, an attempt to bring the success of long CoT to neural machine translation (MT). Specifically, in view of the literature books that might involve similes and metaphors, translating these texts to a target language is very difficult in practice due to cultural differences. In such cases, literal translation often fails to convey the intended meaning effectively. Even for professional human translators, considerable thought must be given to preserving semantics throughout the translation process. To simulate LLMs’ long thought ability in MT, we first mine sentences containing similes or metaphors from existing literature books, and then develop a multi-agent framework to translate these sentences via long thought. In the multi-agent framework, a translator is used to iteratively translate the source sentence under the suggestions provided by an advisor. To ensure the effectiveness of the long thoughts, an evaluator is also employed to judge whether the translation in the current round is better than the previous one or not. In this manner, we collect tens of thousands of long-thought MT data, which is used to train our DRT-o1. The experimental results on literature translation demonstrate the effectiveness of the DRT-o1. Using Qwen2.5-7B and Qwen2.5-14B as the backbones, the improvement brought by DRT-o1 achieves 7.33$\sim$8.26 BLEU and 1.66$\sim$3.36 CometScore. Besides, DRT-o1-7B can outperform QwQ-32B-Preview by 7.82 BLEU and 1.46 CometScore, showing its effectiveness.
@@
Jiaan Wang,  Fandong Mengthanks:   Corresponding author.,  Yunlong Liang,  Jie Zhou Pattern Recognition Center, WeChat AI, Tencent Inc {torchwang,fandongmeng,yunlonliang,withtomzhou}@tencent.com [Uncaptioned image]DRT-o1[Uncaptioned image]DRT-o1-7B[Uncaptioned image]DRT-o1-14B
@@ 1 Introduction
In this paper, we introduce DRT-o1 that aims to bring the success of long thought to neural machine translation (MT). There are two key points in achieving this goal:
@@
Based on the collected long-thought MT samples, we train (SFT) our DRT-o1-7B and DRT-o1-14B using the backbones of Qwen2.5-7B-Instruct and Qwen2.5-14B-Instruct Qwen Team (2024), respectively. Experimental results on literature translation demonstrate the effectiveness of the DRT-o1. For example, DRT-o1-7B outperforms Qwen2.5-7B-Instruct by 8.26 BLEU, 1.31 CometKiwi and 3.36 CometScore. It can also outperform QwQ-32B-Preview by 7.82 BLEU and 1.46 CometScore.
@@ 2.4 Data Statistics
We split the collected 22,264 samples into training, validation and testing sets with 19,264, 1,000 and 2,000 samples, respectively. Table 1 shows the data statistics of DRT-o1 data and previous O1-like data. For Marco-O1 CoT data Zhao et al. (2024), since it is not fully released, we use its demo data to calculate the data statistics. As we can see, the average number of tokens in our synthesized thought reaches 500+ tokens, which is similar to previous math-oriented O1-like CoT data.
@@ 3.1 Experimental Setups
Metrics. Following previous work, we adopt BLEU Papineni et al. (2002), reference-free CometKiwi and reference-based CometScore Rei et al. (2022) to evaluate the model translations. BLEU evaluates n-grams overlap between the generated translations and corresponding references, while CometScore evaluates the semantic similarity of translations against references. CometKiwi uses a language model to judge whether a translation conveys the semantics of the source sentence.
To calculate CometKiwi and CometScore, we leverage the official codes [^2] https://github.com/Unbabel/COMET and the official models [^3] https://huggingface.co/Unbabel/wmt22-cometkiwi-da and https://huggingface.co/Unbabel/wmt22-comet-da. To calculate the BLEU score, we use the sacrebleu toolkit [^4] https://github.com/mjpost/sacrebleu to calulate the corpus-level BLEU.
@@
Backbones. We adopt Qwen2.5-7B-Instruct [^5] https://huggingface.co/Qwen/Qwen2.5-7B-Instruct and Qwen2.5-14B-Instruct [^6] https://huggingface.co/Qwen/Qwen2.5-14B-Instruct Yang et al. (2024) as the backbones of DRT-o1.
Figure 3: An generation case of DRT-o1-14B.
@@
Implementation Details. Llama-Factory Zheng et al. (2024) is used to instruct-tune LLMs. Following Wang et al. (2024), all LLMs are tuned on 8$\times$×NVIDIA A100 GPUs (40G) with 1e-5 learning rate and 32 (8$\times$4) batch size. We use the DeepSpeed optimization Rasley et al. (2020), and set ZeRO-3 optimization. Following Qin et al. (2024), we set the number of training epochs to 3, and the training process costs 70 GPU hours and 124 GPU hours for 7B and 14B models, respectively.
Model BLEU CometKiwi CometScore
Qwen2.5-7B-Instruct 27.02 70.36 76.78
Qwen2.5-14B-Instruct 30.23 72.01 78.84
QwQ-32B-preview 27.46 71.48 78.68
Marco-o1-7B 29.48 71.62 77.41
DRT-o1-7B 35.28 71.67 80.14
DRT-o1-14B 37.56 72.16 80.50
Table 2: Experimental results on literature translation.
@@ 3.2 Main Results
Table 2 shows the results on literature translation. We compare our DRT-o1-7B and DRT-o1-14B with previous Qwen2.5-7B-Instruct, Qwen2.5-14B-Instruct, QwQ-32B-preview and Marco-o1-7B. After instruction tuning on our collected data, DRT-o1-7B outperforms Qwen2.5-7B-Instruct by 8.26 BLEU, 1.31 CometKiwi and 3.36 CometScore. DRT-o1-14B outperforms Qwen2.5-14B-Instruct by 7.33 BLEU, 0.15 CometKiwi and 1.66 CometScore. Besides, DRT-o1-14B achieves the best results in terms of all metrics, showing the effectiveness of long thought in machine translation.
@@
4 Related Work
Recently, O1-like models have shown great performance in reasoning tasks, especially math and coding tasks. After the emergency of OpenAI O1 model OpenAI (2024b), many efforts are given in reproducing OpenAI O1. For example, Qin et al. (2024) propose journey learning, a training paradigm, to encourage LLMs to learn not just shortcuts, but the complete exploration process. With only 327 training samples, the journey learning enhances LLMs’ long-thought ability. Huang et al. (2024) explore the data distillation from existing O1-like models, and show the effectiveness of data distillation. Zhang et al. (2024) leverage Monte Carlo Tree Search (MCTS) to synthesize reasoning-enhanced code data, and train O1-Coder. Marco-o1 Zhao et al. (2024) is proposed to deal with open-ended text generation. Marco-o1 shows the effectiveness of a MCTS-enhanced inference method, and achieves great performance on math problems. Besides, Marco-o1 studies the machine translation cases with long thought, and shows the potentiality of O1-like models in dealing with MT.
5 Conclusion
In this paper, we introduce DRT-o1, an attempt to bring the success of long-thought reasoning to neural machine translation (MT). Specifically, we synthesize the machine translation long-thought samples by a designed multi-agent framework and GPT-4o reformulation. To collect the source sentences that are suitable for translation via long thought, we mine sentences with similes or metaphors from existing literature books. To synthesize the long thought machine translation process for these sentences, a translator, an advisor and an evaluator collaborate to translate the source sentence iteratively. GPT-4o is further employed to enhance the readability and fluency of the thought process collected via the multi-agent framework. Based on the synthesized data, we train DRT-o1-7B and DRT-o1-14B models (using Qwen2.5-7B-Instruct and Qwen2.5-14B-Instruct as backbones), and show their effectiveness in literature translation.

### v2

@@ Abstract
Recently, O1-like models have emerged as representative examples, illustrating the effectiveness of long chain-of-thought (CoT) in reasoning tasks such as math and coding tasks. In this paper, we introduce DRT-o1, an attempt to bring the success of long CoT to neural machine translation (MT). Specifically, in view of the literature books that might involve similes and metaphors, translating these texts to a target language is very difficult in practice due to cultural differences. In such cases, literal translation often fails to convey the intended meaning effectively. Even for professional human translators, considerable thought must be given to preserving semantics throughout the translation process. To simulate LLMs’ long thought ability in MT, we first mine sentences containing similes or metaphors from existing literature books, and then develop a multi-agent framework to translate these sentences via long thought. In the multi-agent framework, a translator is used to iteratively translate the source sentence under the suggestions provided by an advisor. To ensure the effectiveness of the long thoughts, an evaluator is also employed to quantify the translation in each round. In this way, we collect tens of thousands of long-thought MT data, which is used to train our DRT-o1. Using Qwen2.5 and LLama-3.1 as the backbones, DRT-o1 models can learn the thought process during machine translation, and outperform vanilla LLMs as well as existing O1-like LLMs, showing their effectiveness
@@
Jiaan Wang,  Fandong Mengthanks:   Corresponding author.,  Yunlong Liang,  Jie Zhou Pattern Recognition Center, WeChat AI, Tencent Inc {torchwang,fandongmeng,yunlonliang,withtomzhou}@tencent.com [Uncaptioned image]DRT-o1[Uncaptioned image]DRT-o1-7B[Uncaptioned image]DRT-o1-8B[Uncaptioned image]DRT-o1-14B
@@ 1 Introduction
In this paper, our goal is not to achieve competitive performance with OpenAI’s O1 in neural machine translation (MT). Instead, we explore technical routes to bring the success of long thought to MT. To this end, we introduce DRT-o1, a byproduct of our exploration, and we hope it could facilitate the corresponding research in this direction. There are two key points in achieving this goal:
@@
Based on the collected long-thought MT samples, we train (SFT) our DRT-o1-7B, DRT-o1-8B and DRT-o1-14B using the backbones of Qwen2.5-7B-Instruct, Llama-3.1-8B-Instruct Dubey et al. (2024) and Qwen2.5-14B-Instruct Qwen Team (2024), respectively. Experimental results on literature translation demonstrate the effectiveness of the DRT-o1. For example, DRT-o1-14B outperforms vanilla Qwen2.5-14B-Instruct and QwQ-32B-preview in terms of BLEU, CometKiwi, CometScore and GPT-4 evaluators.
@@ 2.3 Long Thought Reformulation
It is also worth noting that during the GPT-4o reformulation, we specify the translation with the highest score $s^{r_{j}}$ as the final translation. Thus, the final translation is not necessarily the last one during refinement, i.e., $t^{r_{n}}$.
Figure 3: The distribution of the number of refinement steps in DRT-o1 data.
@@ 2.4 Data Statistics
We split the collected 22,264 samples into training, validation and testing sets with 19,264, 1,000 and 2,000 samples, respectively. Table 1 shows the data statistics of DRT-o1 data and previous O1-like data. For Marco-O1 CoT data Zhao et al. (2024), since it is not fully released, we use its demo data to calculate the data statistics. [^2] https://github.com/AIDC-AI/Marco-o1 As we can see, the average number of tokens in our synthesized thought reaches 500+ tokens, showing the long thought process in our data.
Figure 3 shows the number of refinement steps/rounds in the DRT-o1 data, which ranges from three to eight. We can find that most samples (73.22%) involve three refinement steps, while only one sample involves eight steps. Furthermore, to provide a deeper understanding of the refinement process, we calculate the average edit distance before and after each refinement step. Specifically, the first three refinement steps cause 21.44, 13.16 and 10.90 character-level edit distance. This observation is consistent with intuition. As the refinement progresses, the magnitude of the modification gradually decreases.
Model reference-free reference-based
GRF CometKiwi GRB BLEU CometScore
Vanilla LLMs
Llama-3.1-8B-Instruct 79.25 70.14 73.30 18.55 74.58
Qwen2.5-7B-Instruct 81.53 70.36 77.92 27.02 76.78
Qwen2.5-14B-Instruct 84.74 72.01 80.85 30.23 78.84
Marco-o1-7B 82.41 71.62 77.50 29.48 77.41
QwQ-32B-preview 86.31 71.48 83.08 27.46 78.68
DRT-o1
DRT-o1-8B (Backbone: ${}_{\text{(Backbone: Llama-3.1-8B-Instruct)}}$ 84.49 70.85 80.80 32.67 78.81
DRT-o1-7B (Backbone: ${}_{\text{(Backbone: Qwen2.5-7B-Instruct)}}$ 85.57 71.78 82.38 35.54 80.19
DRT-o1-14B (Backbone: ${}_{\text{(Backbone: Qwen2.5-14B-Instruct)}}$ 87.19 72.11 83.20 36.46 80.64
Table 2: Experimental results on literature translation.
@@ 3.1 Experimental Setups
Metrics. Following previous work, we adopt BLEU Papineni et al. (2002), reference-free CometKiwi and reference-based CometScore Rei et al. (2022) to evaluate the model translations. BLEU evaluates n-grams overlap between the generated translations and corresponding references, while CometScore evaluates the semantic similarity of translations against references. CometKiwi uses a language model to judge whether a translation conveys the semantics of the source sentence. To calculate CometKiwi and CometScore, we leverage the official codes [^3] https://github.com/Unbabel/COMET and the official models [^4] https://huggingface.co/Unbabel/wmt22-cometkiwi-da and https://huggingface.co/Unbabel/wmt22-comet-da. To calculate the BLEU score, we use the sacrebleu toolkit [^5] https://github.com/mjpost/sacrebleu to calculate the corpus-level BLEU.
In addition, recent studies Kocmi and Federmann (2023); Wang et al. (2023) also show the strong ability of LLMs in NLP evaluation. Thus, we use evaluators implemented using GPT-4o in reference-based and reference-free styles, which we refer to as GRB and GRF, respectively. Since GRB and GRF need the API costs, we randomly select 400 samples from the testing set, and conduct the GRB and GRF evaluation. The evaluation prompts are given in Appendix A.
@@
Backbones. We adopt the following three LLMs as the backbones of our DRT-o1: (1) Llama-3.1-8B-Instruct Dubey et al. (2024) [^6] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct; (2) Qwen2.5-7B-Instruct [^7] https://huggingface.co/Qwen/Qwen2.5-7B-Instruct and (3) Qwen2.5-14B-Instruct Yang et al. (2024) [^8] https://huggingface.co/Qwen/Qwen2.5-14B-Instruct.
@@
Implementation Details. Llama-Factory Zheng et al. (2024) is used to instruct-tune LLMs. Following Wang et al. (2024a), all LLMs are tuned on 8$\times$×NVIDIA A100 GPUs (40G) with 1e-5 learning rate and 8 (8$\times$1) batch size. We use the DeepSpeed optimization Rasley et al. (2020), and set ZeRO-3 optimization. Following Qin et al. (2024), we set the number of training epochs to 3, and the training process costs 70 GPU hours and 124 GPU hours for 7B and 14B models, respectively.
Inference Details. When evaluating model performance on the test set, we use vLLM toolkit Kwon et al. (2023) [^9] https://github.com/vllm-project/vllm to accelerate the model generation. We use the sampling decoding strategy with 0.1 temperature, and set the repetition penalty to 1.05.
Source Sentence Qwen2.5-14B-Instruct QwQ-32B-Preview DRT-o1-14B
Women were weeping and children crying, and all were going as fast as seemingly lay in their power, looking behind now and then as if pursued by some deadly enemy. {CJK}UTF8gbsn 妇女在哭泣,孩子们在哀号,所有人都似乎在尽全力快速前行,不时回头张望,仿佛背后有某个致命的敌人在追赶。 {CJK}UTF8gbsn 妇女们哭泣着,孩子们哭喊着,每个人都在尽力加快脚步,不时回头张望,仿佛背后有致命的敌人在追击。 {CJK}UTF8gbsn 妇女悲泣孩童哀号众人疾奔不时回首,仿佛有致命之敌在后紧追。
Wait through this season of the rain Till suns of autumn dry the plain, Then shall thy giant foe, and all His host and realm, before thee fall. {CJK}UTF8gbsn 等待这雨季过去,直到秋日的阳光晒干原野,那时你的巨敌及其部众和王国都将倒在你面前。 {CJK}UTF8gbsn 忍过雨季漫长夜待秋阳升空晒干大地湿润色,那时,巨敌垮台,他的军队和王国,全在你脚下 {CJK}UTF8gbsn 历尽这雨季,待秋日阳光驱散平原上的最后一丝湿润,那时,你那巨敌,连同他的军队与王国,都将匍匐在你脚下
And if their words had taken another turn…if he himself had been less fastidious about intruding on another man’s secrets…it was cruel to think how thin a film had shut out rescue from all this guilt and misery. {CJK}UTF8gbsn 而如果他们的言辞转向了另一个方向……如果他自己对侵扰他人秘密不那么介意……想到正是那层薄薄的隔阂将救援拒之门外,使得这一切的罪恶和痛苦无法避免,真是残酷。 {CJK}UTF8gbsn 而且,假如他们的谈话走向不同……假如他自己不对干涉他人私事如此谨慎……那时他会意识到,只要一层薄薄的屏障就阻隔了救援,使得所有这些罪恶和痛苦无法解脱,这真是残忍。 {CJK}UTF8gbsn 若他们的言辞稍作转变……若他本人不那么介意侵入他人的心扉……想到这层薄薄的隔阂竟将所有内疚与痛苦隔离于救援之外,实为残忍。
This cold officer upon a monument, who dropped epithets unconcernedly down, would be finer as a dead man, he thought. {CJK}UTF8gbsn 他认为,这个站在纪念碑上的冷漠官员,若死了会更好他不带任何感情地抛下了一些称呼 {CJK}UTF8gbsn 这个冷冰冰的官员站在纪念碑上,毫不在意地抛下一些称号,他想,如果作为一个死人会更出色 {CJK}UTF8gbsn 纪念碑上的冷淡官员,漫不经心地吟咏那些修饰语,他心想,若化为亡者,或许更显尊贵。
I see on your cheek two tears which I know are hot as two sparks, and salt as two crystals of the sea. {CJK}UTF8gbsn 我看见你脸颊上有两滴泪水,我知道它们像两簇火花一样滚烫,像两颗海水结晶一样咸。 {CJK}UTF8gbsn 我见你面颊两颗泪,温热如火花咸涩如海晶 {CJK}UTF8gbsn 我看见你脸颊上的两滴眼泪,它们既如两簇火花般炽热又似两粒海盐般咸涩
The night wore out, and, as he stood upon the bridge listening to the water as it splashed the river-walls of the Island of Paris, where the picturesque confusion of houses and cathedral shone bright in the light of the moon, the day came coldly, looking like a dead face out of the sky. {CJK}UTF8gbsn 夜深了,当他站在桥上,听着水花拍打巴黎岛河岸的声音,月光下房屋与教堂错落有致的景象闪耀着光芒时,冷冰冰的白天来临了,仿佛是从天空中露出的一张死人脸 {CJK}UTF8gbsn 夜晚消逝了,当他站在桥上,聆听水流拍打着巴黎岛的河岸,那里的房屋与大教堂构成了一幅混乱而又美丽的图景,在月光下闪耀着明亮的光芒,而白天则冷酷地到来,仿佛是天空中一张死人脸 {CJK}UTF8gbsn 夜色渐渐消逝,他站在桥上,倾听河水轻拍巴黎城岛的河岸。月光下,房屋与大教堂错落交织闪耀着迷人的光辉。白日冷冰冰地降临,宛如从天际浮现的一张死寂的
Table 3: Case Studies of literature translation. Green indicates good translations, while red indicates bad ones.
3.2 Comparison Models
We use vanilla Llama-3.1-8B-Instruct, Qwen2.5-7B-Instruct and Qwen2.5-14B-Instruct as the comparison models. Besides, two O1-like LLMs, i.e., Marco-o1-7B and QwQ-32B-preview, are also conducted as baselines.
@@ 3.3 Main Results
Table 2 shows the experimental results. After instruction tuning on the long-thought MT training data, DRT-series LLMs significantly outperform the corresponding vanilla backbones. For example, DRT-o1-7B outperforms Qwen2.5-7B-Instruct by 4.04 GRF, 1.42 CometKiwi, 4.46 GRB, 8.52 BLEU and 3.41 CometScore. DRT-o1-8B outperforms Llama-3.1-8B-Instruct by 5.24 GRF, 0.71 CometKiwi, 7.50 GRB, 14.12 BLEU and 4.23 CometScore. Besides, DRT-o1-14B outperforms all baselines (including QwQ-32B-preview and Marco-o1-7B) in terms of all metrics, showing the effectiveness of long thought in MT.
3.4 Case Study
Table 3 provides some literature translation cases of Qwen2.5-14B-Instruct, QwQ-32B-Preview and DRT-o1-14B. With the help of long thought, the translations of DRT-o1-14B align more closely with the conventions of the Chinese language and exhibit a greater literary quality. Besides, in addition to DRT-o1-14B, some translation snippets of other LLMs can also show a great performance (marked in green). This indicates that vanilla LLMs might have the capability to translate literature, and long thought could further activate the capability.
Figure 4: Time cost during inference on the testing set.
3.5 Inference Time Analysis
When evaluating LLMs’ literature translation performance on our test set, we leverage vLLM to accelerate the model generation. Specifically, we use a single NVIDIA A100 GPU (40G) to deploy and call each LLM, and calculate the corresponding generation time in the test set. As shown in Figure 4, the average time costs of DRT-o1 models are significantly higher than vanilla LLMs. This is because DRT-o1 models should first generate the long thought and then provide the final translation, thus needing more inference time. This also indicates that the O1-like LLMs may not be applicable to some scenarios with high real-time requirements.
4 Related Work
4.1 O1-like LLMs
Recently, O1-like LLMs have shown great performance in reasoning tasks, especially math and coding tasks. After the emergency of OpenAI O1 model OpenAI (2024b), many efforts are given in reproducing OpenAI O1. For example, Qin et al. (2024) propose journey learning, a training paradigm, to encourage LLMs to learn not just shortcuts, but the complete exploration process. With only 327 training samples, the journey learning enhances LLMs’ long-thought ability. Huang et al. (2024) explore the data distillation from existing O1-like models, and show the effectiveness of data distillation. Zhang et al. (2024) leverage Monte Carlo Tree Search (MCTS) to synthesize reasoning-enhanced code data, and train O1-Coder. Marco-o1 Zhao et al. (2024) is proposed to deal with open-ended text generation. Marco-o1 shows the effectiveness of a MCTS-enhanced inference method, and achieves great performance on math problems. Besides, Marco-o1 studies the machine translation cases with long thought, and shows the potentiality of O1-like models in dealing with MT.
4.2 Literature Translation
Different from translating standard MT corpora (e.g., news articles), translating literature books is more difficult since it often requires equivalence beyond the word level Thai et al. (2022). Besides, it is also difficult to evaluate literature translation using automatic metrics, and previous literature translation work typically relies on human evaluation Fonteyne et al. (2020); Karpinska and Iyyer (2023). Due to its difficulty, early work is limited to small-scale attempts Genzel et al. (2010); Jones and Irvine (2013); Besacier and Schwartz (2015); Toral et al. (2018). Recently, Karpinska and Iyyer (2023) utilize LLMs to perform literature translation, and show that discourse-level LLM translators achieve better performances compared with sentence-level approaches. Thai et al. (2022) introduce Par3 to benchmark LLMs’ literature translation capability from non-English languages to English. Wang et al. (2024b) propose DelTA, an online document-level MT agent to translate book-level long documents based on short-term memory, long-term memory, summaries and proper noun records.
5 Conclusion
In this paper, we introduce DRT-o1, an attempt to bring the success of long-thought reasoning to neural machine translation (MT). Specifically, we synthesize the machine translation long-thought samples by a designed multi-agent framework and GPT-4o reformulation. To collect the source sentences that are suitable for translation via long thought, we mine sentences with similes or metaphors from existing literature books. To synthesize the long thought machine translation process for these sentences, a translator, an advisor and an evaluator collaborate to translate the source sentence iteratively. GPT-4o is further employed to enhance the readability and fluency of the thought process collected via the multi-agent framework. Based on the synthesized data, we train DRT-o1 models (using Llama-3.1-8B-Instruct, Qwen2.5-7B-Instruct and Qwen2.5-14B-Instruct as backbones), and show their effectiveness in literature translation.
Appendix A GPT-4o Evaluator
For GRB and GRF, we prompt GPT-4o (2024-08-06 version) as the MT evaluator in the reference-based and reference-free manners, respectively. The corresponding prompts borrow from Kocmi and Federmann (2023), and make some adaptions to literature translation (c.f. Figure 5).
Appendix B Complete Generation of DRT-o1